### v1

# Enhancing Person-to-Person Virtual Try-On with Multi-Garment Virtual Try-Off

Abstract
Computer vision is transforming fashion through Virtual Try-On (VTON) and Virtual Try-Off (VTOFF). VTON generates images of a person in a specified garment using a target photo and a standardized garment image, while a more challenging variant, Person-to-Person Virtual Try-On (p2p-VTON), uses a photo of another person wearing the garment. VTOFF, on the other hand, extracts standardized garment images from clothed individuals. We introduce TryOffDiff, a diffusion-based VTOFF model. Built on a latent diffusion framework with SigLIP image conditioning, it effectively captures garment properties like texture, shape, and patterns. TryOffDiff achieves state-of-the-art results on VITON-HD and strong performance on DressCode dataset, covering upper-body, lower-body, and dresses. Enhanced with class-specific embeddings, it pioneers multi-garment VTOFF, the first of its kind. When paired with VTON models, it improves p2p-VTON by minimizing unwanted attribute transfer, such as skin color. Code is available at: https://rizavelioglu.github.io/tryoffdiff/
Figure 1: Virtual try-off results generated by multi-garment TryOffDiff. Rows display the input reference images, model predictions, and the ground truth. Conditioned on the person image and a specified garment class, TryOffDiff reconstructs: upper-body garments VITON-HD (cols 1-3), Dresscode (cols 4-6); lower-body garments (cols 7-9) and dresses (cols 10-12).
1 Introduction
Image-based virtual try-on (VTON) [19] is a key computer vision task aimed at generating images of a person wearing a specified garment. Traditionally, two input images are required: one showing the garment in a standardized form (often from an e-commerce catalog) and another of the target person that needs to be ‘dressed’ [8, 22]. Recent methods focus on a modified formulation where the catalog image is replaced with a photo of another person wearing the target garment, the so-called person-to-person try-on (p2p-VTON). This introduces additional processing complexity [38] as the model does not have access to full garment information.
From an application perspective, VTON offers an interactive shopping experience that helps users make better-informed purchasing decisions. On the research side, it raises intriguing research questions, particularly around human pose detection as well as clothing shape, pattern, and texture analysis [15]. Best-performing models are usually guided generative models focused on creating specific, physically accurate outputs. Unlike general generative tasks that produce diverse outputs, reconstruction requires models to generate images that align with the correct appearance of the garment on a person. However, one major drawback of VTON is the lack of a clearly defined target output, often resulting in stylistic variations that complicate consistent evaluation. As an example, generated images may show garments tucked, untucked, or altered in fit, introducing plausible yet inconsistent visual variations and make it difficult to assess the true quality of the model.
A related task, Virtual Try-OFF (VTOFF) [34], focuses on the reverse goal: generating standardized garment images from photos of clothed persons. Despite sharing challenges with VTON—such as pose analysis, geometric transformations, and preserving textures, patterns, and logos—VTOFF offers a key advantage: its outputs have lower stylistic variability (e.g., no tucked vs. untucked variations), making reconstruction quality easier to assess. However, this line of research remains scarce [37, 32]. From an application perspective, VTOFF can enhance e-commerce by providing consistent garment images for integration into VTON systems, such as OOTDiffusion [40], enabling practical p2p-VTON without ground truth garment images.
In this paper we introduce TryOffDiff, a diffusion-based model tailored to the specific challenges of VTOFF task. The base variant mirrors Stable Diffusion architecture and replaces text conditioning with image conditioning, using SigLIP features [43] processed by an adapter module to capture garment-specific properties such as texture, shape and patterns. Our approach achieves state-of-the-art results on the VITON-HD dataset for upper-body garment reconstruction. To support multi-garment scenarios, we extend TryOffDiff by adding class-specific embeddings to timestep embeddings [24]. We demonstrate our results on the full DressCode dataset and show that the multi-garment TryOffDiff achieves comparable results to garment-specific variants. This makes TryOffDiff the first model to tackle multi-garment virtual try-off. Finally, we integrate TryOffDiff with state-of-the-art VTON models to enable person-to-person try-on. Our results show that this approach outperforms models capable of p2p-VTON by reducing unwanted attribute transfer, such as skin color, from the source person. To summarize, our primary contributions are:
  • TryOffDiff: A diffusion model tuned for virtual try-off, excelling at garment reconstruction.
  • Multi-Garment VTOFF: The first approach capable of handling multiple garment categories, and enabling simultaneous reconstruction of multiple garments, achieved through class-specific embeddings.
  • Enhanced Person-to-Person VTON: Improved performance by feeding TryOffDiff output into VTON systems.
Figure 2: Overview of various fashion image generation pipelines. First row illustrates Virtual Try-On, which synthesizes an image of a person wearing a target garment given input images of person and garment. Second row depicts Virtual Try-Off, which generates an e-commerce-style image of a garment from a photo of a person wearing it. Third row shows Person-to-Person Virtual Try-On, where a garment is transferred from one person to another. Finally, fourth row demonstrates that P2P-VTON can be achieved by integrating VTON and VTOFF models.
2 Related Work
This section reviews prior research relevant to our TryOffDiff model, covering Virtual Try-Off, Image-based Virtual Try-On, and Conditional Diffusion Models. Virtual Try-Off generates standardized garment images from clothed individuals, complementing virtual fitting tasks. Image-based Virtual Try-On synthesizes personalized garment images using catalog or person-worn inputs, supporting advanced variants. Conditional Diffusion Models offer the generative backbone for our approach, enabling controlled synthesis through tailored conditioning mechanisms. These areas highlight the challenges and strategies shaping our work on multi-garment reconstruction.
Virtual Try-Off.
Virtual Try-Off (VTOFF) generates standardized garment images from photos of clothed persons, a task with growing interest due to its applications in e-commerce. Early works, such as TileGAN [42], used a two-stage approach: a U-Net-inspired encoder-decoder to transfer features like color and shape for a coarse image, followed by a pix2pix-based refinement step. With the rise of general-purpose text-to-image diffusion models like Stable Diffusion [28], later works shifted toward text-driven garment generation. ARMANI [47], DiffCloth [48], GarmentAligner [46], and MGD [2] explored finetuning latent diffusion models on garment-caption datasets. However, text-based approaches require highly detailed captions, which are difficult to produce, and often struggle to follow instructions precisely, leading to suboptimal outputs.
To address these limitations, recent approaches adopted image-based conditioning. IGR [31] finetuned Stable Diffusion v1.5 with a dual-tower architecture. Similarly, RAGDiffusion [32] employed SDXL [26] with a dual-tower design and a coarse mask from a complex retrieval-augmented generation. TryOffAnyone [37] finetuned Stable Diffusion v1.5 for VTOFF, training only self-attention layers (similar to CatVTON [10]), without additional image encoders. This resulted in a lightweight yet powerful model, making it a strong baseline. Despite these advances, VTOFF research faces challenges: many studies are not publicly available, limiting open exploration, and most focus solely on upper-body garments, leaving broader garment categories underexplored.
Image-based Virtual Try-On.
The objective of image-based virtual try-on is to create composite images of a target person wearing a specific garment, preserving the person’s identity, pose, and body shape while accurately rendering garment details. CAGAN [19] introduced this task with a cycle-GAN approach, while VITON [15] formalized it as a two-step, supervised framework: warping the garment through non-parametric geometric transformations [4], followed by blending it onto the person. CP-VTON [35] improved warping with a learnable thin-plate spline (TPS) transformation via a geometric matcher, later enhanced by dense flow [16] and appearance flow [14] for better pixel-level alignment of textures and folds. Despite these advanves, warping-based methods struggled with complex garment features like logos and intricate patterns.
Recent efforts shifted to GAN-based and diffusion-based models to overcome these limitations. FW-GAN [12] synthesized try-on videos, while PASTA-GAN [38] adapted StyleGAN2 for person-to-person try-on. However, GANs faced challenges with unstable training and mode collapse, prompting a move toward diffusion models, which offer greater stability. M&M-VTO [49] introduced a single-stage diffusion model for multi-garment try-on using one person image and multiple garment inputs. IDM-VTON [9] used two modules to encode garment semantics, extracting high- and low-level features with cross-attention and self-attention layers. OOTDiffusion [40] leveraged pretrained latent diffusion models, integrating garment features into a denoising UNet via outfitting fusion. In a lighter approach, CatVTON [10] avoided heavy feature extraction, proposing a compact model from a pretrained latent diffusion model with promising results using fewer parameters.
Adapting VTON models for VTOFF is challenging, as VTON often relies on additional inputs like text prompts, keypoints, or segmentation masks, requiring careful adjustment. Importantly, while both VTON and VTOFF tasks involve garment manipulation, they differ fundamentally. VTON models have access to complete garment details, allowing them to primarily focus on warping the item to fit a target pose. In contrast, VTOFF models must work with only partial garment information from a reference image, where occlusions and deformations are common, requiring them to reconstruct missing details from limited visual cues.
Conditional Diffusion Models.
Latent Diffusion Models [28] (LDMs) achieved great success in recent years, providing control over generative process through cross-attention mechanisms [33]. These models support conditioning on various input modalities, including text [5, 3, 13] and images [29, 25, 30]. In text-guided synthesis, ControlNet [44] and T2I-Adapter [23] enhance pretrained models with additional layers for precise spatial control. IP-Adapter [41] further improves flexibility by decoupling cross-attention for text and image features, allowing image-guided generation with optional structural conditions. Prompt-Free Diffusion [39] eliminates text prompts entirely, relying on reference image and optional structural inputs to guide output.
Despite progress, applying these models to garment reconstruction poses challenges. Text-guided methods require highly detailed prompts for each sample to define product attributes, which is impractical. Similarly, existing image-guided models lack the ability to enforce the strict standards of fashion photography [7], limiting their suitability.
3 Methodology
Figure 3: Overview of TryOffDiff. Given a reference image and a class label (e.g. ‘dresses’), the SigLIP [43] image encoder extracts features, which are subsequently processed by a learnable adapter and embedded into the cross-attention layers of the Denoising U-Net. A learnable class label embedding conditions the generation to support multi-garment reconstruction. The model, with trainable Adapter, Embedding, and U-Net components, produces a standardized garment image, which is decoded by a frozen VAE.
This section formally defines the virtual try-off task and outlines TryOffDiff model, including SigLIP-based image conditioning and class-conditioning for multi-garment support.
3.1 Virtual Try-Off
Problem Formulation.
Consider an RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ with height $H\in\mathbb{N}$ and width $W\in\mathbb{N}$, representing a person wearing garments. The VTOFF task is to produce a standardized product image $\mathbf{G}\in\{0,\ldots,255\}^{H\times W\times 3}$, that meets commercial catalog standards. Formally, the goal is to train a generative model that learns the conditional distribution $P(G|C)$, where $G$ and $C$ represent the variables corresponding to garment images and reference images (serving as condition), respectively. Suppose the model approximates this target distribution with $Q(G|C)$. Then, given a specific reference image $\mathbf{I}$ as conditioning input, the objective is for a sample $\hat{\mathbf{G}}\sim Q(G|C=\mathbf{I})$ to resemble a true sample of a garment image $\mathbf{G}\sim P(G|C=\mathbf{I})$ as closely as possible.
3.2 TryOffDiff
The architecture of TryOffDiff mirrors Stable Diffusion [28] (v1.4), a latent diffusion model originally designed for text-to-image generation using CLIP’s [27] text encoder. We replace text prompts for direct image-guided generation.
Image Conditioning.
A core challenge in image-guided generation is effectively incorporating visual features into the conditioning mechanism of the generative model. CLIP’s ViT [27] has become a popular choice for image feature extraction due to its general-purpose capabilities. Recently, SigLIP [43] introduced modifications that improve performance, particularly for tasks requiring more detailed and domain-specific visual representations. Therefore, we use the SigLIP model as image feature extractor and retain the full sequence of token representations in its final layer to preserve spatial information essential for accurate garment reconstruction. These features are processed by a trainable adapter module:
$\mathbf{C}(\mathbf{I})=(\mathrm{LN}\circ\mathrm{Linear}\circ\mathrm{SigLIP})(\mathbf{I})\in\mathbb{R}^{n\times m}$ (1)
where SigLIP embeddings are projected by a linear layer and normalized (LN) [1], following a structure similar to IP-Adapter [41]. The adapted features $\mathbf{C}(\mathbf{I})$ are then incorporated into the denoising U-Net’s cross-attention layers. Specifically, the key $\mathbf{K}$ and value $\mathbf{V}$ of the attention mechanism at each layer are derived from the image features through linear transformations:
$\mathbf{K}=\mathbf{C}(\mathbf{I})\cdot\mathbf{W}_{k}\in\mathbb{R}^{n\times d_{k}},\mathbf{V}=\mathbf{C}(\mathbf{I})\cdot\mathbf{W}_{V}\in\mathbb{R}^{n\times d_{v}}$ (2)
where $\mathbf{W}_{k}\in\mathbb{R}^{m\times d_{k}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{m\times d_{v}}$. This mechanism allows the U-Net to use the reference image’s visual details, improving the precision of the generated garment.
[…]
TryOffAnyone* [37] - - - 72.0 - - 17.2 25.3 - 8.9 21.0
TryOffDiff-single CLIP ViT-B/32 - (257,1024) - 77.9 69.2 46.4 35.3 16.8 7.7 5.2 23.7
CLIP ViT-B/32 Linear+LN (77,768) - 78.8 70.5 46.5 34.3 16.3 7.4 4.7 23.5
SigLIP-B/16 Linear+LN (77,768) - 78.4 69.8 47.3 33.1 16.4 6.6 5.1 22.1
CLIP ViT-B/32 Linear+LN (77,768) 79.8 71.5 47.8 32.0 15.2 5.8 4.1 21.5
SigLIP-B/16 Linear+LN (77,768) 80.3 72.5 49.2 30.1 14.5 5.7 3.9 20.3
Table 1: Training configurations and quantitative results of baseline and ablation studies on VITON-HD-test. The “Pretrain” check mark indicates whether the U-Net is initialized from Stable Diffusion v1.4 (✓) or trained from scratch (-). All ablations mirror the model architecture of Stable Diffusion (v1.4). *Values for TryOffAnyone are taken from the original paper.
Garment Type Conditioning.
Reference images often contain multiple garments, such as a shirt and pants, requiring the model to distinguish which garment to reconstruct. To resolve this ambiguity and enable multi-garment support, we condition the model on text embeddings corresponding to garment categories—“lower body”, “upper body”, or “dress”—aligned with the class labels of the Dress Code dataset. These text embeddings are passed through a linear layer and then combined with timestep embeddings [24] via element-wise addition before being fed into the residual blocks of the diffusion model. This conditioning mechanism guides the denoising process to generate the garment corresponding to specified category.
4 Experiments
Since VTOFF task is newly introduced, few open-source methods exist. We evaluate our proposed TryOffDiff model against the publicly available TryOffAnyone [37] to establish a baseline. To ensure reproducibility, we provide a detailed description of our experimental setup. We adopt DISTS [11] as the primary evaluation metric, supplemented by standard metrics for a comprehensive comparison. Qualitative results are also presented to showcase TryOffDiff’s ability to handle diverse and challenging inputs.
4.1 Experimental Setup
Datasets.
Our experiments are conducted on two publicly available datasets: VITON-HD [8] and Dress Code [22]. Originally designed for VTON, these datasets are also well-suited for VTOFF as they provide image pairs $(\mathbf{I},\mathbf{G})$, where $\mathbf{I}$ is a clothed person and $\mathbf{G}$ is the corresponding garment.
VITON-HD includes $13,679$ high-resolution ($1024\times 768$) image pairs of frontal half-body models with upper-body garments. We preprocessed the dataset by removing duplicates and test set leaks from the training set, resulting in $11,552$ unique pairs for training and $1,990$ for testing.
Dress Code contains $53,792$ high-resolution ($1024\times 768$) person-garment pairs, with $48,392$ pairs for training–comprising dresses ($27,678$), upper-body ($13,563$), and lower-body ($7,151$)–and $5,400$ pairs for testing, evenly split with $1,800$ pairs per category.
For person-to-person try-on, where ground-truth images are unavailable, we randomly pair garments across individuals and provide a text file of input image filenames for reproducibility. For Dress Code, a label (upper-body, lower-body, or dresses) indicates the garment to transfer; this is not needed for VITON-HD, which includes only upper-body garments.
Evaluation Metrics.
To evaluate reconstruction quality, we use established full-reference metrics such as Structural Similarity Index Measure (SSIM) [36] and its multi-scale (MS-SSIM) as well as complex-wavelet (CW-SSIM) variants. To measure perceptual quality, we use common no-reference metrics like Learned Perceptual Image Patch Similarity (LPIPS) [45], Fréchet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [6] metric. To measure the perceptual similarity between images capturing both structural and textural information, we report the Deep Image Structure and Texture Similarity (DISTS) [11] metric. We prioritize the DISTS metric in evaluations because it balances structural and textural information, offering a comprehensive evaluation of image quality.
Implementation Details.
Denoising U-Net weights are initialized from Stable Diffusion v1.4 [28] and subsequently fine-tuned. The Adapter module and class Embedding layer are trained from scratch. The SigLIP encoder, VAE encoder, and VAE decoder are kept frozen throughout the training, which is performed end-to-end. Input reference images are padded along the width to achieve a square aspect ratio and resized to $512\times 512$ to match the pretrained SigLIP and VAE encoder’s format. Garment images undergo the same preprocessing during training. We use SigLIP-B/16-512 as image feature extractor, yielding 1,024 token embeddings of dimension 768, which the adapter–comprising a linear layer and normalization–reduces to $n=77$ conditioning embeddings of dimension $m=768$. For garment conditioning, we embed class labels (e.g. ‘upper-body’, ‘lower-body’, ‘dress’) using a trainable embedding layer. This layer has 3 classes and projects the labels into a 1,280-dimensional space, matching the dimension of the timestep embeddings. These class embeddings are then combined with the timestep embeddings via element-wise addition before being fed into the U-Net’s residual blocks.
Training runs for 200 epochs (approximately 150k iterations) on a single node with four NVIDIA A40 GPUs, taking about 5 days with a batch size of 16. We use the AdamW optimizer [21], with an initial learning rate of 5e-5, which increases linearly from zero over the first 15,000 warmup steps (%10 of total iterations), then decays linearly to zero. A weight decay of 0.01 applies to all parameters except biases and normalization weights. Following [20], we use the Euler noise scheduler (Algorithm 2) with 1,000 steps. Optimization relies on the standard Mean Squared Error (MSE) loss, which measures the distance between the added and the predicted noise at each step [18].
During inference, TryOffDiff uses Euler scheduler with 20 timesteps and a guidance scale of 1.5. On a single NVIDIA A40 GPU, inference on Dress Code test set takes 2.8 seconds per image and requires 10.1GB of memory.
4.2 Ablation Studies
We perform ablation studies to assess key components of TryOffDiff, including the image encoder, adapter module, conditioning feature dimensions, and the impact of finetuning a pretrained U-Net versus training from scratch. The configurations for these experiments are summarized in Tab. 1. In addition, Fig. 4 illustrates the impact of guidance scale and inference steps on FID and DISTS scores, evaluated on the Dress Code-test set using the Euler noise scheduler. This experiment reveals how these hyperparameters influence reconstruction quality, providing insights into optimal inference setting for TryOffDiff.
(a) Guidance Scale
(b) Inference steps
Figure 4: Impact of guidance scale ($s$) and inference steps ($n$) on DISTS and FID scores. Experiments are conducted on DressCode-test with TryOffDiff using the Euler scheduler [20].
4.3 Quantitative Results
We evaluate TryOffDiff’s performance on the VITON-HD and Dress Code datasets, comparing it to baseline TryOffAnyone [37] and category-specific ablations. Results are presented in Tab. 1 (VITON-HD), Tab. 2 (Dress Code), and Tab. 3 (p2p-VTON).
[…]
Dresses
TryOffDiff-single 81.6 76.1 55.6 26.1 18.4 8.2 4.4 20.8
TryOffDiff-multi 81.6 76.1 55.5 26.4 18.9 8.0 4.6 21.2
Upper-body
TryOffDiff-single 80.8 73.8 47.8 31.6 17.1 5.2 4.7 21.6
TryOffDiff-multi 80.2 73.2 48.1 32.3 17.8 5.3 5.5 22.2
Lower-body
TryOffDiff-single 81.1 76.8 55.5 30.0 22.6 9.4 5.9 21.1
TryOffDiff-multi 80.4 75.6 54.1 30.8 22.4 9.0 6.8 21.7
All
TryOffDiff-merged 81.2 75.6 52.9 29.2 11.4 6.7 3.1 21.2
TryOffDiff-multi 78.6 75.4 57.4 25.0 12.8 6.8 4.1 21.3
Table 2: Quantitative comparison of multi-garment VTOFF on DressCode-test. We compare category-specific models (TryOffDiff-single) trained independently on dresses, upper-body, and lower-body garments against a class-conditioned model (TryOffDiff-multi) trained jointly across all garment types. Additionally, we report results for TryOffDiff-merged, which aggregates the outputs of the three category-specific models into a single set. While non-distributional metrics (e.g., SSIM, LPIPS) for merged results reflect simple averages across categories, distributional metrics (e.g., FID, KID) are computed over the merged output set, capturing holistic distributional properties.
VITON-HD Dress Code
[…]
CatVTON 12.0 3.5 3.9 8.4 1.9 3.1
OOTD + GT 10.8 2.8 2.0 7.5 3.4 2.5
OOTD + TryOffDiff 11.9 3.3 2.6 7.9 3.4 2.7
Table 3: Quantitative results for p2p-VTON. We evaluate OOTDiffusion (OOTD) with ground truth (GT) garments and TryOffDiff-predicted garments, alongside CatVTON, a specialized p2p-VTON model. Our TryOffDiff model, when integrated with a VTON model, achieves competitive performance, even though it is not explicitly trained for the p2p task.
The results in Tab. 1 reveal key insights into our design choices. First, finetuning a pretrained U-Net consistently outperforms training from scratch across all metrics, underscoring the advantage of leveraging pretrained weights for initializing the diffusion model. Second, the SigLIP encoder outperforms CLIP, especially in perceptual quality metrics like FID and DISTS, suggesting that SigLIP’s visual representations are more effective for detailed garment reconstruction. Finally, incorporating the adapter module further improves performance by more effectively conditioning the diffusion process with image features.
On VITON-HD dataset (Tab. 1), our method outperforms the baseline across all metrics except LPIPS, where TryOffAnyone achieves a better score, though TryOffDiff excels in global metrics like FID and DISTS which is our primary metric.
In our evaluation on the Dress Code dataset (Tab. 2), we consider: (i) category-specific TryOffDiff models trained individually on dresses, upper-body, and lower-body garments; (ii) a class-conditioned TryOffDiff model trained jointly on all categories using garment-type embeddings; and (iii) an aggregated result, denoted TryOffDiff-merged, where outputs from the three category-specific models are combined post hoc. The class-conditioned model remains competitive across all categories, demonstrating its robustness as a unified solution.
Figure 5: Qualitative comparison of garment-specific and multi-garment TryOffDiff. First row displays the conditioning image. Second row shows dataset-specific reconstructions produced by TryOffDiff model trained on individual datasets: VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Last row shows predictions of multi-garment TryOffDiff trained on full DressCode with garment type conditioning.
For person-to-person virtual try-on (Tab. 3), we integrate TryOffDiff with OOTDiffusion and compare it to CatVTON, a specialized p2p-VTON model. On VITON-HD, OOTD+TryOffDiff performs slightly better than CatVTON. On Dress Code, TryOffDiff, when paired with OOTD, delivers competitive performance, approaching specialized p2p-VTON models like CatVTON, especially in perceptual quality metrics.
These findings highlight TryOffDiff’s versatility and effectiveness, enabling high-quality garment reconstruction for both VTOFF and p2p-VTON tasks. Its improvements in structural similarity and perceptual quality translate into practical benefits, such as more realistic garment rendering in both VTOFF and p2p-VTON applications.
4.4 Qualitative Analysis
Figure 5 illustrates the performance of TryOffDiff in garment reconstruction, comparing two variants: garment-specific models (row 2) and multi-garment model (row 3). The conditioning images, displayed in first row, are sourced from VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Garment-specific models excel in precision, particularly for text and fine details (cols 2 and 5), likely due to the pretrained model initialization. In contrast, the multi-garment model offers versatility, delivering strong performance across diverse garment types and datasets, albeit with slight trade-offs in detail compared to specialized models. Nevertheless, it reconstructs a high level of detail (column 6) and shows robust generalization to VITON-HD, which was unseen during training. In some cases (cols 2, 7 and 9), it even produces better fidelity of color.
Figure 6 illustrates the results of p2p-VTON. Overall, neither method stands out as a definitive winner. CatVTON excels in preserving texture and pattern details but occasionally introduces diffusion artifacts (row 2). It can also transfer unintended attributes, such as skin color, from target model to the source model (row 1)–a limitation not observed in classical try-on models. OOTDiffusion combined with TryOffDiff avoids such attribute transfer issues but may struggle due to masking, resulting in residual source garment attributes in the output (rows 3 and 4).
These results highlight the potential of Virtual Try-Off task and the TryOffDiff model. Although TryOffDiff was not specifically trained for p2p-VTON, its integration with VTON models presents a promising approach, already demonstrating competitive performance compared to state-of-the-art person-to-person virtual try-on methods.
Figure 6: Qualitative comparison for Person-to-Person Virtual Try-On. Columns show: (a) person to be dressed, (b) CatVTON output using a person image with target garment as condition for direct p2p-VTON, (c) OOTDiffusion output with ground truth target garment, and (d) OOTDiffusion output with multi-garment TryOffDiff prediction.
5 Conclusion
In this paper we propose TryOffDiff, a latent diffusion model tailored to Virtual Try-Off. We substitute text conditioning mechanism with adapted SigLIP features to steer the generative process, enabling effective garment reconstructions. TryOffDiff achieves state-of-the art results on VITON-HD. Further experiments were conducted on DressCode, which contains three types of garments: upper-body, lower-body and dresses. While the baseline architecture can be trained separately for each category, we enhance TryOffDiff by introducing class conditioning. This advancement allows a single model to perform multi-garment reconstruction, successfully capturing all garment types with minimal loss in detail quality. To our knowledge, this is the first method to address multi-garment VTOFF in a unified framework.
Moreover, we demonstrate that integrating VTOFF with VTON models enables Person-to-Person Virtual Try-On, achieving performance comparable to specialized models. This pipeline offers practical advantages: it leverages existing VTON models for a wider applications and decouples garment reconstruction from other generative artifacts, simplifying the assessment of p2p-VTON outcomes.
Despite these promising results, challenges persist in preserving complex structures, such as logos and printed designs. Future work could explore advanced generative models, alternative visual conditioning methods and additional losses to enhance detail preservation.
Acknowledgment
This work has been funded by Horizon Europe program under grant agreement 101134447-ENFORCE, and by the German federal state of North Rhine-Westphalia as part of the research funding program KI-Starter. We would like to thank UniZG-FER for providing access to their hardware.